OpenStudy (anonymous):

A package of steaks is $4 more than a package of pork chops. Together they total $18. How much is the package of steaks? (Be sure to type only the price of the package of steaks in the answer blank. Do not use a $ in your answer.) (4 points)

OpenStudy (anonymous):

steaks is 11 dollars

OpenStudy (texaschic101):

s + p = 18 s = p + 4 .... or p = s - 4 now it is just a substitution problem.....can you solve it now ?

OpenStudy (texaschic101):

we are going to sub in s - 4 in for p in the first equation. s + p = 18 s + s - 4 = 18 2s - 4 = 18.....add 4 to both sides 2s = 18 + 4 2s = 22 -- now divide both sides by 2 s = 11.....so there are 11 steaks

OpenStudy (anonymous):

or subtract 4 from 18 which equals 14 then divide by 2 then add 4 to the steaks which is 11-_-

OpenStudy (anonymous):

Y u complex so much

OpenStudy (texaschic101):

it is not complex....it is rather simple. Enough practice and it just comes natural..lol

OpenStudy (anonymous):

s=steaks p=pork p+4=s s+p=18 2p+4=18 -4 -4 2p=14 p=7 s=7+4 A Steak is 11 dollars.

OpenStudy (anonymous):

Whoops, a package of steaks.

OpenStudy (texaschic101):

can't get too many steaks for $11....lol